trent can deliver his newspapers in 30 minutes. It takes Lois 60 minutes to do the same route. How long would it take them to deliver the newspapers if they work together.
Answer by checkley79(3341)

1/30+1/60=1/x
(2+1)/60=1/x
3/60=1/x
3x=60
x=60/3
x=20 minutes working together.
Proof:
1/30+1/60=1/20
3/60=1/20
1/20=1/20